3. Mrs. Mariann Alvarez is approaching a stoplight moving with a


velocity of +25.0 m/s. The light turns yellow, and Ima applies the


brakes and skids to a stop. If Ima's acceleration is -8.00 m/s2, then


determine the displacement of the car during the skidding process.


(Note that the direction of the velocity and the acceleration vectors


are denoted by a + and a - sign.) (3 points)


4. Riza is waiting at a stoplight. When it finally turns green, Riza


accelerated from rest at a rate of a 4.75.00 m/s2 for a time of 4.30


seconds. Determine the displacement of Riza's car during this time


period. (3 points)

1
Expert's answer
2022-04-29T08:27:41-0400

3.


"s=v_0t+\dfrac{at^2}{2}"

"v=v_0+at=0=>t=-\dfrac{v_0}{a}"

Then


"s=v_0(-\\dfrac{v_0}{a})+\\dfrac{a(-\\dfrac{v_0}{a})^2}{2}\t=-\\dfrac{v_0^2}{2a}"

"s=-\dfrac{(25m/s)^2}{2(-8m/s^2)}=39.0625m"

4.


"s=\dfrac{at^2}{2}=\dfrac{4.75m/s^2(4.3s)^2}{2}=43.91375m"